# LLMs as Span Annotators: A Comparative Study of LLMs and Humans

### v3

@@ 1 Introduction
Fine-grained aspects of texts, such as semantic accuracy or coherence, depend on local lexical choices. To reflect these aspects in quality judgments of texts, techniques are needed that provide the appropriate amount of detail. However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect gkatzia2015snapshot; sai2022survey; schmidtova2024automatic. Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models gehrmann2022repairing; liu2023calibrating; wang2023large; gao2024llm.
@@
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale dasanmartino2019finegrained; thomson2020gold; popovic2020informative; kocmi2024error. The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem zheng2023judging; gu2025surveyllmasajudge, allowing task-specific applications kocmi2023gembamqm; hasanain2024large. However, to our knowledge, no study has systematically compared span annotation performance between LLMs and human annotators.
@@
The central focus of our investigation is comparing human annotators and state-of-the-art LLMs on span annotation tasks. We select three span annotation tasks (cf. Section˜3.1): evaluating data-to-text generation thomson2020gold, identifying errors in machine translation kocmi-etal-2024-findings, and detecting propaganda techniques in human-written texts dasanmartino2019finegrained.
@@
    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators (Sections˜3.2 and 5.1).
@@
    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement of qualified crowdworkers (Section˜5.2).
@@
    We discover the sources of model errors by manually analyzing a subset of LLM outputs (Section˜5.3).
@@ 2 Related Work
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts sai2022survey; schmidtova2024automatic. As such, they are unable to quantify more fine-grained aspects (gehrmann2022repairing; 10.1162/tacl_a_00437) and do not correlate well with human judgments novikova-etal-2017-need; reiter-2018-structured. With the emerging LLM-as-a-judge paradigm (gu2025surveyllmasajudge), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (bavaresco2024llms; liu2023geval; sottana2023evaluation; leiter2023eval4nlp; chiang2023can), or free-form feedback (li2024generative; kim2024prometheusa; kim2024prometheus; kartac2025openlgauge). However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases stureborg2024large; koo2023benchmarking; wang2023large.
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (lommel2014multidimensional; mariana2014multidimensional; popovic2020informative; kocmi2024error), where human annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation was applied by thomson2020gold, who introduced a span-based evaluation protocol for annotation of generated basketball match reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection dasanmartino2019finegrained and text summarization subbiah2024reading. Unlike our work, these works focus on span annotation with human annotators.
@@
Early attempts at automating span annotation with ad-hoc guidelines were based on fine-tuned pre-trained encoder models. That includes evaluation of MT (10.1162/tacl_a_00683), D2T generation kasner2021textincontext or text summarization goyal2022snac, as well as propaganda detection martino2020semeval2020; goffredo2023argumentbased; piskorski2023multilingual. Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on recent work that applies LLMs as a task-specific evaluation tool (kocmi2023gembamqm; fernandes-etal-2023-devil; hasanain2024large; kasner2024traditional; chang2024booookscore; zouhar2024aiassisted; kartac2025openlgauge; ramponi2025finegrained). Our work is the first that unifies diverse span annotation tasks within a single framework and systematically compares the performance of LLMs to human annotators.
@@ 3.2 Span Annotation with LLMs
To obtain the annotations, we follow the setup of kasner2024traditional: we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^2] Following castillo2024structured, we ensure that the reason field is generated first.
@@ 3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. We consider the following similarity metrics:
@@
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by dasanmartino2019finegrained. These measures are on matching annotations, adjusted to give partial credit to imperfect matches:
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow dasanmartino2019finegrained and hasanain2024large in using the $\gamma$ score (mathet2015unified) as a complementary metric. The metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from -$\infty$ to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (krippendorff1980content), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux2021.
@@ 4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs sharma2022innovations; celikyilmaz2021evaluation. The inputs $X$ are the structured data used to generate the output text $Y$.
We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks balloccu2024leak; dong2024generalization; jiang2024investigating. Instead of using an existing dataset, we use the Quintd tool kasner2024traditional to download structured inputs from multiple public APIs. [^3] We selected two of the existing domains: openweather for generating weather forecasts and gsmarena for generating phone descriptions. We also add the football domain (using RapidAPI - API-Football) for generating soccer game reports. To obtain output texts for the structured data, we prompt LLMs in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. See Appendix B.1 for more details.
To gather annotations for our dataset, we use crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task tseng2020best; iskender2020best; huang2023incorporating; zhang-etal-2023-needle. Our process of collecting annotations proceeded in two stages, following the setup of zhang-etal-2023-needle: (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations. See Appendix B.2 for more details on collecting annotations.
@@ 4.1.2 MT-Eval: Identifying Errors in Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (kocmi2024findings). The system outputs were annotated with the Error Span Annotation (ESA) protocol (kocmi2024error) by professional translators.
@@
The original dataset has nearly 50k model outputs, making it too extensive for our evaluation campaign. Therefore, we used a balanced sample of all available system outputs for 10 sampled inputs for each pair of languages, making up 2,854 examples in total.
@@ 4.1.3 Propaganda: Propaganda Technique Detection
For the Propaganda task, we use the dataset of dasanmartino2019finegrained containing news collected mostly from on-line propagandistic sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Inputs $X$ are empty for this task, as all annotated categories are intrinsic to the evaluated text $Y$.
@@ 4.2 Collecting LLM annotations
  • instruction-tuned models: Llama 3.3 70B grattafiori2024llama, GPT-4o hurst2024gpt, and Claude 3.7 Sonnet anthropic2025claude,
  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer marjanovic2025deepseek. DeepSeek-R1 distilled 70B version guo2025deepseek, o3-mini openai2025o3mini, and Gemini 2.0 Flash Thinking google2025geminithinking.
@@ 5.1 Prompting Techniques
Figure 2: Comparison of LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. For $F_{1}$ score, the shadow bar denotes its soft variant. The dashed horizontal lines for D2T-Eval denote agreement between our human annotators. Detailed results are included in Tables˜11, 12, 13 and 14.
@@
Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1. This observation is aligned with guo2025deepseek, who note that few-shot prompting degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for further experiments.
@@
DeepSeek-R1 generally outperforms the same-sized Llama 3.3. Its superiority is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o. A notable exception to this trend is the non-reasoning Claude 3.7 Sonnet, which scores mostly on par with o3-mini and excels at MT-Eval.
@@
For D2T-Eval, we compare model results with an average IAA on a subset of examples annotated by two human annotators. Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of dasanmartino2019finegrained, who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher than LLMs (the best LLM score being $\gamma=0.16$ for Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter property is reflected in the large difference between the soft and hard $F_{1}$ scores. [^5] We omit the comparison with human IAA in MT-Eval. While the WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, these examples take up only a small fraction and exhibit high variance between language pairs.
@@ 5.3 Manual Analysis of LLM Annotations
To gain more insights into the qualitative aspects of LLM annotations, we manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. [^7] We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. Without access to the annotation source, we classified the annotations and their explanations as Correct, Partially correct, Wrong category, Incorrect, and Undecidable.
@@ 6 Discussion
The IAA between LLMs and human annotators is only moderate, suggesting LLMs cannot straightforwardly replace human annotators. However, using LLMs may be a reasonable option in scenarios such as D2T-Eval based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves. In other cases, when deciding whether to employ LLMs as span annotators, one needs to balance desired output quality with other practical aspects. Here, LLMs provide better flexibility, shorter response times, and lower costs. One should also consider the quality of available human annotators, as even qualified crowdworkers make similar amounts of errors as LLMs. It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks veselovsky2023artificial. A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations zouhar2024aiassisted.

### v4

@@ 1 Introduction
Fine-grained aspects of texts, such as semantic accuracy or coherence, depend on local lexical choices. To reflect these aspects in quality judgments of texts, techniques are needed that provide the appropriate amount of detail. However, most automatic evaluation metrics for Natural Language Generation (NLG) assign only singular scores for the whole text per each evaluated aspect Gkatzia and Mahamood (2015); Sai et al. (2023); Schmidtová et al. (2024). Although numerical values make it easy to rank systems, these metrics are too crude and susceptible to biases or miscalibration of the underlying models Gehrmann et al. (2023); Liu et al. (2024); Wang et al. (2024); Gao et al. (2024).
@@
Despite its advantages, span annotation has not yet been widely applied in automatic NLG evaluation. The method traditionally required human annotators, making it costly and difficult to scale Da San Martino et al. (2019); Thomson and Reiter (2020); Popovic (2020); Kocmi et al. (2024c). The LLM-as-a-judge paradigm recently emerged as a promising solution to this problem Zheng et al. (2023); Gu et al. (2024), allowing task-specific applications Kocmi and Federmann (2023); Hasanain et al. (2024). However, to our knowledge, no study has systematically compared span annotation performance between LLMs and human annotators.
@@
The central focus of our investigation is comparing human annotators and state-of-the-art LLMs on span annotation tasks. We select three span annotation tasks (cf. Section˜3.1): evaluating data-to-text generation Thomson and Reiter (2020), identifying errors in machine translation Kocmi et al. (2024a), and detecting propaganda techniques in human-written texts Da San Martino et al. (2019).
@@
    We establish that with structured outputs and detailed annotation guidelines, LLMs can serve as robust span annotators, yielding relevant spans for all three annotation tasks we work with (Sections˜3.2 and 5.1).
@@
    We show that LLMs have moderate inter-annotator agreement with human annotators overall, but can reach the agreement level among verified crowdworkers who passed a qualification task (Section˜5.2).
@@
    We discover the sources of model errors by manually analyzing a subset of LLM annotation outputs (Section˜5.3).
@@ 2 Related Work
Automatic NLG metrics traditionally assess text quality by measuring similarity to human-written reference texts Sai et al. (2023); Schmidtová et al. (2024). As such, they are unable to quantify more fine-grained aspects (Gehrmann et al., 2023; Freitag et al., 2021) and do not correlate well with human judgments Novikova et al. (2017); Reiter (2018). With the emerging LLM-as-a-judge paradigm (Gu et al., 2024), LLMs have been applied as evaluators on various tasks, using simple numeric scoring (Bavaresco et al., 2025; Liu et al., 2023; Sottana et al., 2023; Leiter et al., 2023; Chiang and Lee, 2023), or free-form feedback (Li et al., 2024; Kim et al., 2024a, b; Kartáč et al., 2025). However, as these outputs are not firmly grounded in text, they tend to miss fine-grained aspects and are influenced by LLM biases Stureborg et al. (2024); Koo et al. (2024); Wang et al. (2024).
@@
In machine translation (MT), span annotation is a long-standing component of protocols such as MQM or ESA (Lommel et al., 2014; Mariana, 2014; Popovic, 2020; Kocmi et al., 2024c), where human annotators mark erroneous spans in translations. In data-to-text (D2T) generation, span annotation was applied by Thomson and Reiter (2020), who introduced a span-based evaluation protocol for annotation of generated basketball match reports. Span annotation is also used to judge intrinsic text qualities, such as coherence or use of rhetorical devices, in tasks such as propaganda detection Da San Martino et al. (2019) and text summarization Subbiah et al. (2024). Unlike our work, these works focus on span annotation with human annotators.
@@
Early attempts at automating span annotation with ad-hoc guidelines were based on fine-tuned pre-trained encoder models. That includes evaluation of MT (Guerreiro et al., 2024), D2T generation Kasner et al. (2021) or text summarization Goyal et al. (2022), as well as propaganda detection Martino et al. (2020); Goffredo et al. (2023); Piskorski et al. (2023). Automating span annotation with LLMs is more flexible and benefits from increasing LLM capabilities. We build on work that applies LLMs as a task-specific evaluation tool (Kocmi and Federmann, 2023; Fernandes et al., 2023; Hasanain et al., 2024; Kasner and Dušek, 2024; Chang et al., 2024; Zouhar et al., 2025; Kartáč et al., 2025; Ramponi et al., 2025). Furthermore, Semin et al. (2026) recently investigated various strategies for automatic span annotation with LLMs. Our work is the first that systematically compares the performance of LLMs to human annotators.
@@ 3.2 Span Annotation with LLMs
To obtain the annotations, we follow the setup of Kasner and Dušek (2024): we request the list of annotations in JSON format, using constrained decoding with a fixed JSON scheme to ensure that the output is syntactically valid. We require each annotation to contain the fields reason (the explanation $r_{i}$), text (the textual content of the span), and type (the integer index of the error category $c_{i}$). [^2] Following Castillo (2024), we ensure that the reason field is generated first.
@@ 3.3 Evaluating Span Annotations
To compare annotations automatically, we need a notion of similarity between two sets of annotations $\mathcal{A}=\{A_{1},A_{2},\ldots,A_{|Y|}\}$ and $\hat{\mathcal{A}}=\{\hat{A}_{1},\hat{A}_{2},\ldots,\hat{A}_{|Y|}\}$ over a set of texts $\mathcal{Y}=\{Y_{1},Y_{2},\ldots,Y_{|Y|}\}$. Note that basic inter-annotator agreement metrics such as Cohen’s $\kappa$ (Cohen, 1960) are not applicable in our case, as they require a fixed set of annotation units, while the number and position of spans in the span annotation task may differ Mathet et al. (2015). Therefore, we consider the following similarity metrics:
@@
To quantify the degree of alignment between individual annotations, we compute precision, recall, and $F_{1}$ as defined by Da San Martino et al. (2019). These measures are on matching annotations, adjusted to give partial credit to imperfect matches:
@@
The $F_{1}$ score is sensitive to varying span granularities and does not consider near-matches with no overlap or agreement by chance. To this end, we follow Da San Martino et al. (2019) and Hasanain et al. (2024) in using the $\gamma$ score (Mathet et al., 2015) as a complementary metric. The metric builds the best possible alignment between the sets of annotations $A_{Y}$ and $\hat{A}_{Y}$ and computes the “disorder” of this alignment based on the positional and categorical dissimilarities of aligned annotations. The score ranges from -$\infty$ to 1, where 1 is achieved when the annotations are perfectly aligned. The $\gamma$ score extends Krippendorff’s $\alpha$ (Krippendorff, 1980), another popular metric, by computing the category-aware span alignments. We use the implementation of Titeux and Riad (2021).
@@ 4.1.1 D2T-Eval: Evaluation of Data-to-text Generation
In D2T-Eval, we use span annotation to evaluate semantic accuracy and stylistic aspects of data-to-text generation outputs Sharma et al. (2022); Celikyilmaz et al. (2020). The inputs $X$ are the structured data used to generate the output text $Y$.
We use D2T-Eval as a control task to mitigate the effects of data contamination: the fact that the performance of the model might be inflated by previous exposure to publicly available benchmarks Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024). Instead of using an existing dataset, we use the Quintd tool Kasner and Dušek (2024) to download structured inputs from multiple public APIs. [^3] We selected two of the existing domains: openweather for generating weather forecasts and gsmarena for generating phone descriptions. We also add the football domain (using RapidAPI - API-Football) for generating soccer game reports. To obtain output texts for the structured data, we prompt LLMs in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. Note that we do not need to deal with the factuality of outputs here, as the sole purpose of the texts is being the input to the annotation process (in fact, having some number of errors is desirable). See Appendix B.1 for more details.
To gather annotations for our dataset, we use crowdworkers from Prolific.com. We apply best practices for gathering human annotations, including an iterative process to refine annotation guidelines and preselecting the best-performing annotators using a qualification task Tseng et al. (2020); Iskender et al. (2020); Huang et al. (2023); Zhang et al. (2023). Our process of collecting annotations proceeded in two stages, following the setup of Zhang et al. (2023): (1) a qualification task for pre-selecting skilled annotators, and (2) the main task for collecting the annotations. See Appendix B.2 for more details on collecting annotations.
@@ 4.1.2 MT-Eval: Identifying Errors in Machine Translation
For MT-Eval, we use the dataset of system outputs from the WMT 2024 general shared task (Kocmi et al., 2024b). The system outputs were annotated with the Error Span Annotation (ESA) protocol (Kocmi et al., 2024c) by professional translators.
@@
The original dataset has nearly 50k model outputs, making it too extensive for our evaluation campaign. Therefore, we used a balanced subsample: For each of the nine language pairs, we randomly sample ten input translation segments. We then take all available system outputs for these 90 input segments, making up 2,854 examples in total.
@@ 4.1.3 Propaganda: Propaganda Technique Detection
For the Propaganda task, we use the dataset of Da San Martino et al. (2019) containing news collected mostly from on-line propagandistic sources. The token-level annotations in the dataset created by expert annotators cover 18 categories of logical fallacies and persuasion techniques. We use the test split for our experiments. Inputs $X$ are empty for this task, as all annotated categories are intrinsic to the evaluated text $Y$.
@@ 4.2 Collecting LLM annotations
  • instruction-tuned models: Llama 3.3 70B Grattafiori et al. (2024), GPT-4o Hurst et al. (2024), and Claude 3.7 Sonnet Anthropic (2025),
  • reasoning models: [^4] By reasoning models we understand the models that use extra inference time to generate a thinking trace before providing the answer Marjanović et al. (2025). DeepSeek-R1 70B DeepSeek-AI (2025), o3-mini OpenAI (2025), and Gemini 2.0 Flash Thinking Deepmind (2025).
@@ 5.1 Prompting Techniques
Figure 2: Comparison between LLMs using $\mathcal{P}_{\texttt{base}}$ and human annotators. Rows represent different tasks (see Section 4.1), columns show different annotation similarity metrics (see Section 3.3). For the $F_{1}$ score, the shadow bar denotes its soft variant. The dashed horizontal lines denote agreement between our human annotators for D2T-Eval (the agreement is not available for the external datasets). More detailed results are included in Tables˜11, 12, 13 and 14.
@@
Few-shot prompting ($\mathcal{P}_{\texttt{5shot}}$) brings ambivalent results, increasing Llama 3.3 scores but doing the opposite for DeepSeek-R1. This observation is aligned with DeepSeek-AI (2025), who note that few-shot prompting degrades the performance of DeepSeek-R1. Given these considerations, we decided to use $\mathcal{P}_{\texttt{base}}$ for further experiments.
@@
DeepSeek-R1 generally outperforms the same-sized Llama 3.3.555The models are comparable as the 70B distilled variant of DeepSeek-R1 is based on Llama 3.3 70B. See Section A.1 for details. Its superiority is most pronounced on D2T-Eval ($F_{1}$-score of 0.23 vs. 0.18, $\gamma$ score of 0.19 vs. 0.11). The same observation applies to OpenAI models, where o3-mini outperforms GPT-4o. A notable exception to this trend is the non-reasoning Claude 3.7 Sonnet, which scores mostly on par with o3-mini and excels at MT-Eval.
@@
For D2T-Eval, we compare model results with an average IAA on a subset of examples annotated by two human annotators. Here, o3-mini, Claude 3.7 and Gemini 2.0 mostly reach or surpass human agreement. For Propaganda, the upper bound of IAA is the result of Da San Martino et al. (2019), who report $\gamma=0.31$ for annotators before consolidation. This score is substantially higher than LLMs (the best LLM score being $\gamma=0.16$ for Claude 3.7 Sonnet). However, this task used expert annotators and has the largest number of categories. The latter property is reflected in the large difference between the soft and hard $F_{1}$ scores. [^6] We omit the comparison with human IAA in MT-Eval. While the WMT24 dataset for MT-Eval contains examples annotated with a pair of annotators, these examples take up only a small fraction and exhibit high variance between language pairs.
@@ 5.3 Manual Analysis of LLM Annotations
To gain more insights into the qualitative aspects of LLM annotations, we manually analyzed the quality of LLM annotations on 216 samples from D2T-Eval and Propaganda. [^8] The analysis was split among 7 authors of this paper. While we did not do double annotation due to lack of time, we discussed any unclear cases throughout the process. We do not include MT-Eval in the manual analysis due to our insufficient expertise in most target languages. For each model, we sampled three annotations per category in D2T-Eval and one annotation per category in Propaganda. Without access to the annotation source, we classified the annotations and their explanations as Correct, Partially correct, Wrong category, Incorrect, and Undecidable.
@@ 6 Discussion
The IAA between LLMs and human annotators is only moderate, suggesting LLMs cannot straightforwardly replace human annotators. However, using LLMs may be a reasonable option in scenarios based on crowdworkers, where the strongest LLMs reach the average IAA between human annotators themselves. In other cases, when deciding whether to employ LLMs as span annotators, one needs to balance desired output quality with other practical aspects. Here, LLMs provide better flexibility, shorter response times, and lower costs. One should also consider the quality of available human annotators, as even qualified crowdworkers (i.e., those who passed a qualification task) make similar amounts of errors as LLMs. It can be also assumed that LLM-based span annotation will benefit from future increases in LLM capabilities, while crowdworkers may increasingly rely on LLM to complete tasks Veselovsky et al. (2023). A promising solution seems to be a hybrid approach in which LLMs pre-annotate the text and humans post-edit the annotations Zouhar et al. (2025).
@@
Is the task meaningful despite the low scores?
As pointed out by an anonymous reviewer of this paper, the low annotation accuracy – as found by our manual analysis – may indicate a fundamental limitation of the proposed evaluation setup. Span annotation is indeed complex and leaves more room for subjectivity than more straightforward annotation such as simple labels or scores. However, we argue that the detailed actionable feedback gained through span annotation outweighs the increased noise, both in terms of explainability of the outputs and their potential for further processing.